### v1

@@ 2 Preliminaries
2.1 Stable matchings, priority violations, lattices, EADAM, and rotations
Stable matchings and priority violations.
We study matchings between applicants and institutions (outside of the exposition in Section 1, where we referred to students and schools). Denote the set of applicants by $\mathcal{A}$ and institutions by $\mathcal{I}$. For simplicity we assume that all matchings are one-to-one, i.e. that institutions have capacity $1$. We typically denote matchings by $\mu,\nu,\xi$, etc., applicants by $d$ (mnemonic: doctor), and institutions by $h$ (mnemonic: hospital); let $\mu(d)$ denote the match of $d$ in $\mu$ (possibly equal to $\emptyset$ if $d$ is unmatched). Each agent has strict preferences over the other side. We denote the preferences of applicant $d$ as $\succ_{d}$, and write $h_{1}\succ_{d}h_{2}$ if $d$ prefers institution $h_{1}$ to institution $h_{2}$. Applicants may have partial preference lists, indicating that they prefer to remain unmatched over being matched to any institution not on their list. The “preferences” of the institutions, also denoted by $\succ_{h}$ for an institution $h$, are referred to as priorities. This reflects the fact that we think of the institutions as “objects to be consumed” without preferences of their own.555 For example, in student-school matching, priorities of schools might be determined by pre-specified policies based on a student’s socio-economic status or distance from the school. For simplicity, we assume that each institution ranks all applicants in their priority lists.
We think of stability in terms of avoiding violating applicants’ priorities, as follows:
Definition 2.1.
@@ 2.1 Stable matchings, priority violations, lattices, EADAM, and rotations
For an applicant $d$, institution $h$, and matching $\mu$, we say that $d$’s priority at $h$ is violated in $\mu$ if $h\succ_{d}\mu(d)$ and $d\succ_{h}\mu(h)$. We say that $d$’s priority is violated in $\mu$ if there exists such an $h$. We denote by $\operatorname*{vio}(\mu)\subseteq\mathcal{A}$ the set of applicants whose priority is violated in $\mu$.
A matching $\mu$ is stable if no applicant’s priority is violated in $\mu$, i.e., if $\operatorname*{vio}(\mu)=\emptyset$.
The lattice of stable matchings.
@@
A lattice under a partial order $\leq$ is a set $Q$ such that every pair of elements in $Q$ have a greatest lower bound and a least upper bound in $Q$. The greatest lower bound of $x,y\in Q$, also called their meet in $Q$ and denoted $x\wedge_{Q}y$, is the element $z\in Q$ such that $z\leq x$ and $z\leq y$, and moreover for each $w\in Q$ such that $w\leq x$ and $w\leq y$, we have $w\leq z$. Likewise, the least upper bound of $x,y\in Q$, also called their join in $Q$ and denoted $x\vee_{Q}y$, is the element $z\in Q$ such that $z\geq x$ and $z\geq y$, and for each $w\in Q$ such that $w\geq x$ and $w\geq y$, we have $w\geq z$. Since we study different lattices on sets of matchings, we often emphasize the relevant lattice $Q$ using the notation $\wedge_{Q}$ and $\vee_{Q}$. A lattice $Q$ is distributive if, for each $x,y,z\in Q$, we have $\mu\wedge_{Q}(\nu\vee_{Q}\xi)=(\mu\wedge_{Q}\nu)\vee_{Q}(\mu\wedge_{Q}\xi)$.
We now explain the classical lattice structure of stable matchings. We say that matching $\mu$ (weakly) dominates matching $\nu$, denoted $\mu\geq\nu$, if for each applicant $d$ we have $\mu(d)\succeq_{d}\nu(d)$. We write $\mu>\nu$ if $\mu\geq\nu$ and $\mu\neq\nu$. For two matchings $\mu,\nu$, let the coordinatewise maximum of $\mu$ and $\nu$ be the function $\xi$ from applicants to institutions given by $\xi(d)=\max_{\succ_{d}}\{\mu(d),\nu(d)\}$. Analogously, let their coordinatewise minimum of $\mu$ and $\nu$ be such that $\xi(d)=\min_{\succ_{d}}\{\mu(d),\nu(d)\}$. Note that for general $\mu$ and $\nu$, the coordinatewise maximum/minimum may not be a matching.
Theorem 2.2 (See, e.g., [GI89]).
@@
The set of stable matchings is always nonempty, and forms a lattice under the dominance ordering. In this lattice, the least upper bound of any two stable matchings is their student-by-student maximum, and the greatest lower bound is analogously their student-by-student minimum. Moreover, a lattice $Q$ is isomorphic to a lattice of stable matchings for some set of preferences and priorities if and only if $Q$ is distributive.
We refer to the applicant-optimal stable matching as the applicant-proposing deferred acceptance (denoted $\mathsf{APDA}$) matching. Likewise, we term the institution-optimal version institutions-proposing deferred acceptance (denoted $\mathsf{IPDA}$).
Pareto improving stable matchings and EADAM.
A matching $\mu$ is Pareto-optimal if there does not exist a $\nu\neq\mu$ with $\nu\geq\mu$. The literature on Pareto-improving stable matchings has identified a central mechanism which is Pareto-optimal and dominates $\mathsf{DA}$, namely, [Kes10]’s efficiency-adjusted DA mechanism ($\mathsf{EADAM}$). One approach to defining $\mathsf{EADAM}$, pioneered by [TY14], is as follows:
Definition 2.3.
For a profile of preferences and priorities $P$, define the Kesten-Tang-Yu sequence as follows. Let $\mu_{1}$ denote the outcome of $\mathsf{APDA}(P)$ under preference and priorities $P=P_{1}$. Call an institution $h$ under-demanded in round $i$ if $h$ never rejects any applicant during the run of $\mathsf{APDA}(P_{i})$ (additionally, the “institution” $\emptyset$ representing going unmatched is always considered under-demanded). For each $i\geq 2$, let $P_{i}$ denote modifying the profile of preferences as follows: for each student $d$ matched in $\mu_{i-1}$ to an institution $h$ which is under-demanded in round $i-1$, remove all institution strictly before $h$ in $d$’s preference list. There must exist an $N$ such that $\mu_{n}=\mu_{m}$ for each $n,m>N$. The outcome of efficiency-adjusted DA ($\mathsf{EADAM}$) is $\mu_{N}$.
[TY14] prove that at least one institution is under-demanded in each round, and thus $\mathsf{EADAM}$ can be computed via at most $N$ runs of $\mathsf{APDA}$, where $N$ is the number of institutions. Moreover, $\mathsf{EADAM}$ is Pareto-optimal and dominates all stable matchings.
$\mathsf{EADAM}$ is closely related to priority-neutral matchings and will feature into our analysis. The same is true of another prior notion, namely the set of legal matchings [EM20], but we defer defining legal matchings and their properties (e.g., from [FZ22]) until they are needed in Section 4.
Rotations.
@@
Throughout the paper, we use the notion of rotations [GI89] to describe different matchings. A rotation is an ordered list of agents $\alpha=[(h_{1},d_{1}),\allowbreak(h_{2},d_{2}),\allowbreak\ldots,\allowbreak(h_{k},d_{k})]$, where each $h_{i}\in\mathcal{I}$ and $d_{i}\in\mathcal{A}$. If $\mu$ is a matching where $d_{i}$ is paired to $h_{i}$ for each $i=1,\ldots,k$, then we let $\alpha\mu$ denote the matching which differs from $\mu$ only by matching $d_{i}$ to $h_{i+1}$ for $i=1,\ldots,k$ modulo $k$. (The action of a rotation can be visualized as moving the parentheses in $\alpha$ over one place, modulo $k$.) Matching $\alpha\mu$ is also called the elimination of rotation $\alpha$ from $\mu$. In Section 3, we use rotations solely as notation, i.e., one can treat them as names for different matching. In Section 4, we will exploit many properties of rotations from the literature as needed; see [GI89] or [EIV23] for a treatment.666 Note that we will use the convention that $\alpha\mu\geq\mu$, i.e. all applicants weakly prefer $\alpha\mu$ to $\mu$. In contrast to this, [GI89] (resp., [EIV23]) define rotations such that $\mu\geq\alpha\mu$, i.e., such that all men (resp., workers) weakly prefer $\mu$ to $\alpha\mu$. Other than that this difference in convention, the notions are identical. We write $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}\mu$ to mean $\alpha_{1}(\alpha_{2}(\ldots(\alpha_{\ell}\mu)\ldots))$; when the “base” matching $\mu$ is understood, we also simply denote this by $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}$.
2.2 Priority-neutral matchings and their join and meet operations
@@
We now turn to the central notion of this paper, priority-neutral matchings [Ren22]. In words, a matching $\mu$ is priority-neutral if it’s impossible to make any student whose priority is violated in $\mu$ better off, without violating the priority of some student who is made worse off. To fully spell out this definition, we introduce the terminology “priority-correcting adjustment” to aid in discussing matchings which are not priority-neutral. Namely, we have:
Definition 2.4.
For matchings $\xi$ and $\nu$, we say that $\nu$ is a priority-correcting adjustment of $\xi$ if:
  1. 1.
@@
A matching $\mu$ is priority-neutral if there does not exist any priority-correcting adjustment of $\mu$. We denote the set of priority-neutral matchings by $\mathsf{PN}$.
$\mathsf{PN}$ generalizes the set of stable matchings, and [Ren22] proves that $\mathsf{PN}$ also forms a lattice. Moreover, [Ren22] proves that the set $\mathsf{PN}$ is closed under the coordinatewise maximum operation, so the join operation in the $\mathsf{PN}$ lattice is also given by the coordinatewise maximum. Formally:
@@
For any profile of priorities and preferences, $\mathsf{PN}$ is a lattice under $\leq$, and this lattice contains every stable matching. Additionally, the student-pessimal element of $\mathsf{PN}$ is the institution-optimal stable matching, and the student-optimal element of $\mathsf{PN}$ equals the outcome of $\mathsf{EADAM}$. Finally, for any two matchings $\mu,\nu$ in $\mathsf{PN}$, the least upper bound of $\mu$ and $\nu$ in $\mathsf{PN}$ equals their coordinatewise maximum.
A simple example showing greatest lower bounds need not be the coordinatewise minimum.
Perhaps surprisingly, we will show in this paper that the greatest lower bound between two matchings in $\mathsf{PN}$ may not be their coordinatewise minimum. Before proceeding, we first describe how this type of situation is possible using a simple example.
Example 2.6.
Consider the following three matchings:
$\mu_{0}$ =$=\{(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{3}),(h_{4},d_{4}),(h_{5},d_{5}),(h_{6},d_{6})\}$
$\mu_{L}$ =$=\{(h_{1},d_{2}),(h_{2},d_{1}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{5}),(h_{6},d_{6})\}$
$\mu_{R}$ =$=\{(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{6}),(h_{6},d_{5})\}$
In words, relative to $\mu_{0}$, matching $\mu_{L}$ separately swaps the matches of $d_{1},d_{2}$ and (separately) of $d_{3},d_{4}$, and $\mu_{R}$ swaps $d_{3},d_{4}$ and $d_{5},d_{6}$. Let the applicant’s preferences be such $\mu_{L}$ and $\mu_{R}$ are both preferred to $\mu_{0}$ by every applicant who changes match between them. Then, within the set of matchings $Q=\{\mu_{0},\mu_{L},\mu_{R}\}$, matchings $\mu_{L}$ and $\mu_{R}$ have a greatest lower bound $\mu_{L}\wedge_{Q}\mu_{R}=\mu_{0}$, which is different from their coordinatewise minimum $[(h_{1},d_{1}),(h_{2},d_{2}),(h_{3},d_{4}),(h_{4},d_{3}),(h_{5},d_{5}),(h_{6},d_{6})]$.
Notation for joins and meets.
In Section 3, we embed an example like 2.6 into $\mathsf{PN}$. In view of such examples, we introduce separate notation for the coordinatewise minimum and the greatest lower bound of two matchings in $\mathsf{PN}$. Namely, let $\mu\wedge_{\mathsf{L}}\nu$ denote the coordinatewise minimum of $\mu$ and $\nu$,777 This notation is inspired by the legal set of matchings $\mathsf{L}$ from [EM20]; as we will exploit in Section 4, coordinatewise minimums $\wedge_{\mathsf{L}}$ give greatest lower bounds in the legal lattice $\mathsf{L}$. and let $\mu\wedge_{\mathsf{PN}}\nu$ denote their greatest lower bound in $\mathsf{PN}$. Since by Theorem 2.5, least upper bounds equal coordinatewise maximums in $\mathsf{PN}$ (and since this is also true in all other matching lattices we consider), we use the notation $\mu\vee\nu$ for least upper bounds (i.e., this always denotes the coordinatewise maximum).

### v2

@@ 2 Preliminaries
2.1 The Stable Matching Lattice
We begin by recalling the definition of stable matchings and their essential properties.
Stabile Matchings and Priority Violations.
We study one-to-one matchings between applicants $\mathcal{A}$ and institutions $\mathcal{I}$ (outside of some exposition in the introduction, where we referred to students and schools, respectively). We typically use notations like $\mu,\nu,\xi$ for matchings and (to avoid the common variables $a$ and $i$) we write $d\in\mathcal{A},h\in\mathcal{I}$ (mnemonic: doctors and hospitals). For $d\in\mathcal{A}$, we write $\mu(d)\in\mathcal{I}\cup\{\emptyset\}$ for $d$’s match, where $\mu(d)=\emptyset$ denotes going unmatched (and likewise with $\mu(h)$ for $h\in\mathcal{I}$).
Agents on each side have a strict ordinal preference list over the other side, i.e., rankings $\succ_{d}$ over $\mathcal{I}$ for each $d\in\mathcal{A}$, and likewise $\succ_{h}$ over $\mathcal{A}$ for $h\in\mathcal{I}$. We refer to these rankings of applicants as their preferences, but by convention, we refer to the rankings of institutions as priorities (reflecting the fact that, as in many school choice mechanisms, we think of the priorities as fixed and/or determined by policies, instead of reflecting the institution’s preferences). Following [Ren22], we assume for simplicity that each institution ranks all applicants; we allow applicants’ lists to be partial (i.e., they can rank some “outside option” $\emptyset$, which represents going unmatched, above some institutions).
Definition 2.1 (Stable matchings and priority violations).
@@ 2.1 The Stable Matching Lattice
A matching $\mu$ is unstable if there exists a $d\in\mathcal{A}$ and $h\in\mathcal{I}$ such that $h\succ_{d}\mu(d)$ and $d\succ_{h}\mu(h)$. In this case, we say that $d$’s priority is violated (at $h$). If there are no such $d$ and $h$, we say that $\mu$ is stable. Let $\mathsf{Stab}$ denote the set of all stable matchings.
For a matching $\mu$, denote the set of applicants whose priority is violated in $\mu$ by
$\operatorname*{vio}(\mu)=\{d\in\mathcal{A}\ \mid\ \exists h\in\mathcal{I}:\text{ $d\succ_{h}\mu(h)$ and $h\succ_{d}\mu(d)$}\}.\qed$
We let $\mathsf{APDA}$ denote the applicant-optimal stable matching, i.e. the outcome of the applicant-proposing deferred acceptance algorithm. Likewise, let $\mathsf{IPDA}$ denote the institution-optimal / institution-proposing matching. We typically consider a fixed profile of applicant preferences and institution priorities $P$, and thus suppress the dependence on $P$ in our notation $\mathsf{APDA}$, $\mathsf{IPDA}$, and other matching mechanism. When we need to refer to $P$ explicitly, we denote the matchings $\mathsf{APDA}(P)$ and $\mathsf{IPDA}(P)$.
(Stable) Lattices.
@@
This paper studies lattice structures on sets of matchings. A (finite) lattice $Q$ is a partial order $\leq$ over a (finite) set $Q$ such that every pair of elements in $Q$ have a meet and a join. The meet (also called the greatest lower bound) of $x,y\in Q$, denoted $x\wedge_{Q}y$, is the element $z\in Q$ such that $z\leq x$ and $z\leq y$, and moreover for each $w\in Q$ such that $w\leq x$ and $w\leq y$, we have $w\leq z$. Likewise, the join (also called the least upper bound) of $x,y\in Q$, denoted $x\vee_{Q}y$, is the element $z\in Q$ such that $z\geq x$ and $z\geq y$, and for each $w\in Q$ such that $w\geq x$ and $w\geq y$, we have $w\geq z$. A lattice $Q$ is distributive if, for each $x,y,z\in Q$, we have $\mu\wedge_{Q}(\nu\vee_{Q}\xi)=(\mu\wedge_{Q}\nu)\vee_{Q}(\mu\wedge_{Q}\xi)$.
We say that matching $\mu$ (weakly) dominates matching $\nu$, denoted $\mu\geq\nu$, if for each applicant $d$ we have $\mu(d)\succeq_{d}\nu(d)$. We write $\mu>\nu$ if $\mu\geq\nu$ and $\mu\neq\nu$. Throughout this entire paper, the matchings are ordered by dominance.
For two matchings $\mu,\nu$, let the coordinatewise maximum (resp., minimum) of $\mu$ and $\nu$ be the function $\xi$ from applicants to institutions given by $\xi(d)=\max_{\succ_{d}}\{\mu(d),\nu(d)\}$ (resp., $\xi(d)=\min_{\succ_{d}}\{\mu(d),\nu(d)\}$). In some (but not all) of the latices we consider, joins/meets are given by coordinatewise maxima/minima.
It is classically known that stable matchings form a distributive lattice with joins/meets given by coordinatewise maxima/minima, and that all distributive lattices arise as the set of stable matchings [Knu76, Bla84]. For exposition, see [GI89, EIV23].
Theorem 2.2 (Distributive Lattices Characterize Stable Matchings).
@@
The set of stable matchings is always nonempty, and forms a lattice under the dominance ordering. In this lattice, the join (resp., meet) of any two stable matchings is their coordinatewise maximum (resp., minimum). Moreover, a lattice $Q$ is isomorphic to a lattice of stable matchings for some set of preferences and priorities if and only if $Q$ is distributive.
Denoting Matchings Using Rotations.
The distributive lattice of stable matchings additionally has a compact and useful representation in terms of a concept called “rotations” [GI89, EIV23]. We recall the full details of this representation in Section 4.1. For ease of exposition, we begin by using rotations only as notation for denoting different matchings, as follows.
@@
A rotation is an ordered list of agents $\alpha=[(h_{1},d_{1}),\allowbreak(h_{2},d_{2}),\allowbreak\ldots,\allowbreak(h_{k},d_{k})]$, where each $h_{i}\in\mathcal{I}$ and $d_{i}\in\mathcal{A}$. If matching $\mu$ has $\mu(h_{i})=d_{i}$ for each $i=1,\ldots,k$, then we let $\alpha\mu$, which is called the elimination of rotation $\alpha$ from $\mu$, denote the matching which differs from $\mu$ only by matching $d_{i}$ to $h_{i+1}$ for $i=1,\ldots,k$ modulo $k$. Hence, the action of a rotation can be visualized as moving the parentheses in $\alpha$ over one place, modulo $k$. Moreover, we will always consider rotations $\alpha$ eliminated at matchings $\mu$ such that all applicants weakly prefer $\alpha\mu$ to $\mu$; this convention allows us to think about matchings “higher” in the lattice as corresponding to “larger” sets of rotations, which we find intuitive. [^7] In contrast to this, [GI89] (resp., [EIV23]) define rotations such that $\mu\geq\alpha\mu$, i.e., such that all men (resp., workers) weakly prefer $\mu$ to $\alpha\mu$. Other than this change in convention, the notions are identical. We adopt our convention—that larger sets of rotations correspond to matchings preferred more by the applicants—in an attempt to make working with priority-neutral matchings intuitive, since in this setting one side of the market (the applicants) are distinguished, and we seek to improve their welfare. We write $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}\mu$ to mean $\alpha_{1}(\alpha_{2}(\ldots(\alpha_{\ell}\mu)\ldots))$; when the “base” matching $\mu$ is understood, we also simply denote this by $\alpha_{1}\alpha_{2}\ldots\alpha_{\ell}$.
2.2 The Priority-Neutral Matching Lattice
As discussed in the introduction, there are important reasons to look beyond the set of stable matchings. We first discuss a specific matching mechanism $\mathsf{EADAM}$ which improves applicants’ matches beyond their stable ones, and will be important for our analysis. Then we recall the main notion we are interested in: priority-neutrality.
Pareto-Optimality and $\mathsf{EADAM}$.
When $\mu$ dominates $\nu$, i.e., $\mu>\nu$, we also say that $\mu$ Pareto-improves $\nu$. Matching $\mu$ is Pareto-optimal if there does not exist a $\nu$ with $\nu>\mu$. The literature on Pareto-improving stable matchings has identified a central mechanism which is Pareto-optimal and dominates $\mathsf{DA}$, namely, [Kes10]’s efficiency-adjusted DA mechanism ($\mathsf{EADAM}$). One approach to defining $\mathsf{EADAM}$, pioneered by [TY14], is as follows:
Definition 2.3 ($\mathsf{EADAM}$ and the Kesten-Tang-Yu sequence).
For a profile of preferences and priorities $P$, define the Kesten-Tang-Yu sequence as follows. Let $\mu_{1}$ denote the outcome of $\mathsf{APDA}(P)$ under preference and priorities $P=P_{1}$. Call an institution $h$ under-demanded in round $i$ if $h$ never rejects any applicant during the run of $\mathsf{APDA}(P_{i})$ (additionally, the “institution” $\emptyset$ representing going unmatched is always considered under-demanded). For each $i\geq 2$, let $P_{i}$ denote modifying the profile of preferences as follows: for each applicant $d$ matched in $\mu_{i-1}$ to an institution $h$ which is under-demanded in round i1i-1, remove all institutions strictly before $h$ in $d$’s preference list. Then, let $\mu_{i}=\mathsf{APDA}(P_{i})$. There must exist an $N$ such that $\mu_{n}=\mu_{N}$ for each $n\geq N$. The outcome of efficiency-adjusted DA ($\mathsf{EADAM}$) is $\mu_{N}$. ∎
[TY14] prove that at least one institution is under-demanded in each round, and thus $\mathsf{EADAM}$ can be computed via at most $N$ runs of $\mathsf{APDA}$, where $N$ is the number of institutions. Moreover, $\mathsf{EADAM}$ is Pareto-optimal and dominates all stable matchings. $\mathsf{EADAM}$ is closely related to our main definition of interest, priority-neutrality, and will feature into our analysis.
@@ 2.2 The Priority-Neutral Matching Lattice
We now turn to the main object of study in this paper: the lattice of priority-neutral matchings [Ren22]. In words, a matching $\mu$ is priority-neutral if it’s impossible to make any applicant whose priority is violated in $\mu$ better off, without violating the priority of some applicant who is made worse off. To fully spell out this definition, we introduce the terminology “priority-correcting adjustment” to aid in discussing matchings which are not priority-neutral. Namely, we have:
Definition 2.4 (Priority-Neutrality).
Say that matching $\nu$ is a priority-correcting adjustment (abbreviated PCA) of matching $\xi$ if:
  1. (1)
@@
A matching $\mu$ is priority-neutral if there does not exist any priority-correcting adjustment of $\mu$. We denote the set of priority-neutral matchings by $\mathsf{PN}$. ∎
Priority-correcting adjustments capture failures of priority-neutrality; indeed, [Ren22] argues they “correct” some applicant $d$’s priority violation (without unduly harming others). In other words, priority-neutral matchings allow for some applicants’ priority to be violated, but only when such applicants cannot be improved without harming others and also violating their priority.
Observe that, since stable matchings violate no applicant’s priority, every stable matching is priority-neutral. Thus, $\mathsf{PN}$ generalizes the set of stable matchings, and [Ren22] proves that $\mathsf{PN}$ also forms a lattice. Moreover, [Ren22] proves that the set $\mathsf{PN}$ is closed under the coordinatewise maximum operation, and identifies $\mathsf{IPDA}$ and $\mathsf{EADAM}$ as the applicant-pessimal and applicant-optimal elements of $\mathsf{PN}$, respectively. Formally:
@@
For any profile of priorities and preferences, $\mathsf{PN}$ is a lattice under $\leq$, and this lattice contains every stable matching. Additionally, the applicant-pessimal element of $\mathsf{PN}$ is $\mathsf{IPDA}$, and the applicant-optimal element of $\mathsf{PN}$ is $\mathsf{EADAM}$. Finally, for any two matchings $\mu,\nu$ in $\mathsf{PN}$, the least upper bound of $\mu$ and $\nu$ in $\mathsf{PN}$ equals their coordinatewise maximum.
Note, however, that [Ren22] leaves as an open question whether the greatest lower bound of two matchings in $\mathsf{PN}$ is their coordinatewise minimum. Our first result below resolves this question.